# Rest energy effects for non-relativistic Fermi fields in external potentials

### v3

@@ IV Discussion
The non-relativistic counterpart of the Dirac equation should not be considered necessarily as its non-relativistic limit. There are external potentials in which the rest energy contribution is not trivial. We can still change the reference point and shift the energy levels by a constant value. However, the rest energy effects are not erased by the shift. With the rest energy included, the GLL equation leads to a physical picture some elements of which cannot be seen in the framework of the LL equation. We demonstrated this for a non-relativistic Fermi field in a spherical well type potential of finite depth.
In this potential, the spin effects do not show themselves, and the results are the same for spin up and spin down components. This allows us to see clearly the rest energy effects. One of the new elements is the fact that the existence of bound states depends on the rest energy of the Fermi particle. The particle entering the region of the well is confined to it if the rest energy differs from zero and exceeds a minimum value. The minimum is determined by the depth and width of the well. The larger the rest energy, the more bound states we get. Otherwise, the particle is not ”captured” by the well. If it is in a bound state and its rest energy goes below the minimum value, this destroys the bound state, and the particle leaves the well.

### v4

@@ 4 Discussion
The non-relativistic counterpart of the Dirac equation should not be considered necessarily as its non-relativistic limit. The GLL equation gives us a new insight into the physical consequences of rest energy. There are external potentials in which the rest energy contribution becomes non-trivial. We can still change the reference point and shift the energy levels by a constant value. However, the rest energy effects are not erased by the shift. With the rest energy included, the GLL equation leads to a physical picture some elements of which cannot be seen in the framework of the LL equation. We have demonstrated this for a non-relativistic Fermi field in a spherical well type potential of finite depth.
In this potential, the spin effects do not show themselves, and the results are the same for spin up and spin down components. This allows us to see clearly the rest energy effects. One of the new elements is the fact that the existence of bound states and their number depend not only on the depth and width of the well, as it was known before, but also on the rest energy of the Fermi particle. The particle entering the region of the well is confined to it if the rest energy differs from zero and exceeds a minimum value. The minimum is determined by the depth and width of the well. Otherwise, the particle is not captured by the well. If it is in a bound state and its rest energy goes below the minimum value, this destroys the bound state, and the particle leaves the well.
The finite well potentials have multiple applications in condensed matter systems. The spherical well type potentials are used in semiconductor quantum dots. The experimental realization of potentials with the rest energy contribution would provide its impact on bandgap structure and functions of quantum dots. If observed experimentally, the threshold value ${\mu}_{\rm min}$ could be used to estimate the value of the rest mass $m_{0}$.
The approach presented can be applied to a more general choice of ${\varphi}$ in Eq.(36). Then the solution of the GLL equation will exhibit both the rest energy and spin effects including spin-orbit coupling.